OpenStudy (anonymous):

Reduce the following rational expression to lowest terms, if possible. (Problem in comments)

OpenStudy (anonymous):

lol whats the problem ?

OpenStudy (anonymous):

\[\frac{ 20a+60 }{ 5a+15 }\]

OpenStudy (anonymous):

I have to factor the top by 10 & the bottom by 5, right?

hartnn (hartnn):

top by 20 :)

OpenStudy (anonymous):

lol idk! i hate math lol this would not be a problem ik the answer to @hartnn might be able to help

OpenStudy (anonymous):

ohh lol he's here. nvm

OpenStudy (anonymous):

lmaoo of course he's here lol I thought 20, but wanted to do 10 for some reason lol ok \[\frac{ 20(a+3) }{ 5(a+3) }\]

hartnn (hartnn):

thats correct what gets cancelled ?

hartnn (hartnn):

\(\huge \dfrac{ \cancel{20}^4\cancel{(a+3)} }{ \cancel5\cancel{(a+3)} }\)

OpenStudy (anonymous):

4!

hartnn (hartnn):

good! :)

hartnn (hartnn):

oh, and you also need restrictions, right ?

OpenStudy (anonymous):

No, no restrictions lol

hartnn (hartnn):

x cannot take the value when \(\large denominator =0\) so go for 5a+15 =0 even if you don't need, its better to know :)